You bought a stock three months ago for $43.53 per share. The stock paid no dividends. The current share price is $47.24 What is the APR of your investment? APR = 34.09 % What is the EAR % of your investment?
Answer #1

Purchase Price = $43.53
Selling Price = $47.24

3-month Return = (Selling Price - Purchase Price) / Purchase Price
3-month Return = ($47.24 - $43.53) / $43.53
3-month Return = 0.085225 or 8.5225%

EAR = (1 + 3-month Return)^4 - 1
EAR = (1 + 0.085225)^4 - 1
EAR = 1.3870 - 1
EAR = 0.3870 or 38.70%